OpenStudy (anonymous):

If 1/a +1/b = 10, what is the value of a+b ?

OpenStudy (anonymous):

Choices are : 1/10 , 2/5 , 1 , 1/10 or can't be determined.

OpenStudy (anonymous):

i believe they would be 2/5

OpenStudy (anonymous):

i assume a=b because their is only one answer not one for a and one for b?

OpenStudy (anonymous):

its 2/5. I am sure

OpenStudy (anonymous):

yup i agree

OpenStudy (anonymous):

\[\frac{1}{a}+\frac{1}{b}=10\]Combine fractions on the LHS.\[\frac{a+b}{a b}=10\]\[a+b=10a b\]

OpenStudy (anonymous):

"can't be determined"